OpenStudy (anonymous):

Find the surface area of a cylinder with a base area of 49 π in2 and a height that is twice the radius of its base.

OpenStudy (anonymous):

pi*r^2 = 49 equate to get r r = sqrt of 49/pi height = r*2 substitute r and height to the surface area cylinder formula: 2*pi*r^2+2*pi*r*h

OpenStudy (anonymous):

I got 32.3pi^2. I'm pretty sure thats wrong though . . .